# MegaSaM: Accurate, Fast, and Robust Structure and Motion from Casual Dynamic Videos

### v1

@@ 1 Introduction
Extracting camera parameters and scene geometry from a set of images is a fundamental problem in computer vision, commonly referred to as structure from motion (SfM) or Simultaneous Localization and Mapping (SLAM). While decades of research have yielded mature algorithms for stationary scenes with large camera baselines, these methods often falter when applied to casual monocular videos captured in uncontrolled setting [26, 79]. Such videos, frequently captured by handheld cameras, typically exhibit limited camera motion parallax (e.g., nearly stationary or rotational cameras) and a broad range of focal lengths, and often include moving objects and scene dynamics. Recent efforts to address these challenges have focused on two primary strategies: optimizing camera and scene geometry by fine-tuning mono-depth networks or reconstructing radiance fields [36, 78, 79, 35]; or combining intermediate estimates derived from monocular video, such as depth, flows, long-term trajectory and motion segmentation, into a global optimization framework [26, 6, 77, 80, 35]. However, these approaches are computational expensive or brittle when applied to unconstrained videos that feature long time duration, unconstrained camera paths, or complex scene dynamics.
@@
In this work, we present MegaSaM, a full pipeline for accurate, fast and robust camera tracking and depth estimation from in-the-wild monocular videos of dynamic scenes. Our approach combines the strengths of several prior works, leading to results with previously unachievable quality, as shown in Fig. 1. In particular, we reexamine and extend prior deep visual SLAM framework for camera tracking. A defining feature of deep visual SLAM systems like DROID-SLAM [60] is that they adopt a differentiable bundle adjustment (BA) layer that iteratively updates scene geometry and camera pose variables, and intermediate predictions are learned from large amounts of data through camera and flow supervision. We find that such a learned layer is critical for achieving accurate and efficient camera pose estimation for the more challenging case of dynamic video. Building on this foundation, one of our key innovations for handling dynamic scene is to integrate monocular depth priors and motion probability maps into a differentiable SLAM paradigm.
@@ 2 Related Work
SLAM and SfM are used to estimate camera parameters and 3D scene structure from video sequences or unstructured image collections. Conventional approaches tackle this problem by first estimating 2D correspondence between images through feature matching [56, 1, 52, 45, 46, 5, 38, 8, 25, 57] or photometric alignment [39, 9, 10]. They then optimize 3D point locations and camera parameters by minimizing reprojection or photo-consistency errors through bundle adjustment (BA) [62].
Recently, deep visual SLAM and SfM systems have emerged that adopt deep neural networks to estimate pairwise or long-term correspondences [2, 58, 66, 74, 61, 60, 64, 55, 7, 18, 21, 19], to reconstruct radiance fields [34, 11, 42] or global 3D point clouds [28, 67]. While these methods demonstrate accurate camera tracking and reconstruction, they typically assume predominantly static scenes and sufficient camera baselines between frames. Therefore, their performance can degrade significantly or fail entirely in the presence of scene dynamics or limited camera parallax.
Several recent works share similar goals with ours in addressing these limitations. Robust-CVD [29] and CasualSAM [79] jointly estimate camera parameters and dense depth maps from dynamic videos by optimizing a spatially varying spline or fine-tuning monocular depth networks. Particle-SfM [80] and LEAP-VO [6] first infer moving object masks based on long-range trajectories, then use this information to downweight the contribution of the features during bundle adjustment. Concurrent work, MonST3R [77], adopts a 3D point cloud representation from DuST3R [67] and localizes cameras via an additional alignment optimization. Our approach shares similar ideas, but we show that performance can be significantly improved by coupling the differentiable SLAM system with the intermediate predictions of underlying dynamic scenes.
@@
Recent work on monocular depth prediction has shown strong generalization on in-the-wild single images by training deep neural networks on large amounts of synthetic and real-world data [30, 31, 48, 75, 47, 50, 72, 73, 76, 44, 22, 13]. However, these single-image models tend to produce temporally inconsistent depth from videos. To overcome this issue, prior techniques propose to fine-tune mono-depth models by performing test-time optimization [36, 78] or to use transformer or diffusion models to directly predict video depths [70, 54, 20]. Our approach follows the spirit of the first paradigm, but we show that we can achieve better video depth quality without resorting to expensive network fine-tuning for every video.
@@
Several recent works have adopted time-varying radiance field representations to perform dynamic scene reconstruction and novel view synthesis from in-the-wild videos [40, 41, 32, 33, 12, 68, 27, 65, 35, 71]. Our work is orthogonal to most of these techniques since most radiance field reconstruction methods require camera parameters or video depth maps as inputs, and our outputs can be used as inputs to these systems.
@@ 3 MegaSaM
Given an unconstrained, continuous video sequence $\mathcal{V}=\{I_{i}\in\mathcal{R}^{H\times W}\}^{N}_{i=1}$ our goal is to estimate camera poses $\hat{\mathbf{G}}_{i}\in SE(3)$, focal length $\hat{f}$ (if unknown), and dense video depth maps $\mathcal{D}=\{\hat{D}_{i}\}^{N}_{i=1}$. Our approach does not pose any constraints on the camera and object motions present in the input video. Our camera tracking and video depth estimation modules build upon prior deep visual SLAM (in particular, DROID-SLAM [60]) and casual structure and motion [79] frameworks, respectively.
@@ 3.1 Deep visual SLAM formulation
Deep visual SLAM systems like DROID-SLAM [60] are characterized by a differentiable, learned bundle adjustment (BA) layer that iteratively updates structure and motion parameters. In particular, they keep track of two state variables while processing a video: a per-frame low-resolution disparity map $\hat{\mathbf{d}}_{i}\in\mathcal{R}^{\frac{H}{8}\times\frac{W}{8}}$, and camera poses $\hat{\mathbf{G}}_{i}\in SE(3)$. These variables are updated iteratively during both training and inference stages through the differentiable BA layer, which operates over a set of image pairs from a frame-graph $(I_{i},I_{j})\in\mathcal{P}$, built dynamically to connect frames with overlapping field of view.
@@
Differentiable bundle adjustment. DROID-SLAM assumes known focal length, but focal length is typically not known a priori for in-the-wild videos. Therefore, we optimize camera poses, focal length and disparity by iteratively minimizing a weighted reprojection cost between the current flows predicted by the network and the rigid-motion ones derived from camera parameters and disparity [16]:
@@
where $\Delta\boldsymbol{\xi}=(\Delta\mathbf{G},\Delta\mathbf{d},\Delta f)^{T}$ is the parameter updates of the state variables, $\mathbf{J}$ is the Jacobian of reprojection residuals w.r.t. the parameters, and $\mathbf{W}$ is a diagonal matrix containing $\hat{\mathbf{w}}_{ij}$ from each frame pair. $\lambda$ is a damping factor predicted by the network during each BA iteration. We can separate camera parameters (including poses and focal length) and disparity variables by dividing the approximated Hessian on the LHS of Eq. 3 into following block matrix form:
@@
Since only a single disparity variable is included in each pairwise reprojection term from Eq. 2, $\mathbf{H}_{\mathbf{d}}$ in Eq. 4 is a diagonal matrix, and hence we can efficiently compute parameter updates using the Schur complement trick [62], which leads to fully differentiable BA update:
@@
Training. The flow and uncertainty predictions are trained end-to-end from a collection of synthetic video sequences of static scenes:
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
I Image I Motion probability
@@ 3.2.1 Training
Disparity and camera initialization. DROID-SLAM initializes disparity $\hat{\mathbf{d}}$ by simply setting it to a constant value of $1$. However, we find that this initialization fails to perform accurate camera tracking on videos with limited camera baselines and complex scene dynamics. Inspired by recent work [32, 65, 33, 35], during both the training and inference stages, we perform data-driven initialization by integrating a mono-depth prior. During training, we initialize $\hat{\mathbf{d}}$ with disparity from DepthAnything [72] with the estimated global scale and shift borrowed from the ground truth depth of each training sequence. For each training sequence, we initialize first two camera poses to the ground-truth to remove the gauge ambiguity, and initialize the camera focal length by randomly perturbing the ground truth value by $25\%$.
@@ 3.2.2 Inference
Initialization and frontend tracking. Similar to training, we incorporate both mono-depth and focal length predictions into the inference pipeline. In particular, we initialize per-frame disparity maps $\hat{\mathbf{d}}_{i}$ with metric aligned monocular disparity $D^{\text{align}}_{i}=\hat{\alpha}D^{\text{rel}}_{i}+\hat{\beta}$, where $D^{\text{rel}}_{i}$ is per-frame affine-invariant disparity from [72], and per-video global scale and shift parameters $(\hat{\alpha},\hat{\beta})$ are estimated through median aligning $D^{\text{rel}}_{i}$ with an additional metric depth estimates $D^{\text{abs}}_{i}$ from UniDepth [44]: $\hat{\alpha}_{i}=\frac{D^{\text{abs}}_{i}-\text{median}_{i}(D^{\text{abs}}_{i})}{D^{\text{rel}}_{i}-\text{median}(D^{\text{rel}}_{i})}$; $\hat{\beta}=\text{median}\left(D^{\text{abs}}_{i}-\hat{\alpha}D^{\text{rel}}_{i}\right)$. [^1] We combine two mono-depth models since they have complementary advantages: DepthAnything provides more accurate and consistent depths, whereas UniDepth provides scene scale and shift estimates. The UniDepth model also predicts the focal length of each frame; we use the median estimate across the video frames to obtain an initial focal length estimate $\hat{f}$, which is fixed within the frontend stage.
@@
To see why, we explore the approximate Hessian matrix from the linear system in Eq. 4. As shown by Goli et al. [14], given the posterior, $p(\theta|\mathcal{I})$, we can use Laplace approximations to estimate the covariance $\Sigma$ of the variables through the inverse Hessian: $\Sigma_{\theta}=-\mathbf{H}(\theta^{*})^{-1}$, where $\theta^{*}$ is the MAP estimate of the parameters and $\Sigma_{\theta}$ denotes the epistemic uncertainty of estimated variables [23]. Since inverting the full Hessian is computationally expensive when the number of input frames is large, we follow Ritter et al. [49] and approximate $\Sigma_{\theta}$ through the diagonal of the Hessian:
@@ 3.3 Consistent depth optimization
In particular, we follow CasualSAM [79] and perform an additional first-order optimization on video depths along with per-frame aleatoric uncertainty maps. Our objective consist of three cost functions:
@@
where $\mathcal{C}_{\text{flow}}$ denotes pairwise 2D flow reprojection loss, $\mathcal{C}_{\text{temp}}$ is temporal depth consistency loss, and $\mathcal{C}_{\text{prior}}$ is scale invariant mono-depth prior loss. We derive 2D optical flow at the original frame resolution from an off-the-shelf module [59].
Note that our design has a few differences compared to CasualSAM: (i) instead of performing a time-consuming mono-depth network fine-tuning, we construct and optimize a sequence of variables for disparity and uncertainty over the input video; (ii) we fix camera parameters instead of jointly optimizing cameras and depths during optimization; (iii) we adopt surface normal consistency and multi-scale depth gradient matching losses [30, 51] to replace the depth prior loss used in CasualSAM [79]. We find that these modifications lead to much faster optimization times as well as more accurate video depth estimates. We refer readers see the supplementary material for more details of our losses and optimization scheme.
@@ 4 Experiments
CasualSAM [79] 0.036 0.013 0.20 0.067 0.019 0.47 1.6m
LEAP-VO [6] 0.041 0.023 0.17 - - - 1.3s
ACE-Zero [3] 0.053 0.028 0.30 0.065 0.028 1.92 10s
Particle-SfM [80] 0.062 0.032 1.26 0.057 0.038 1.64 21s
RoDynRF [35] 0.110 0.049 1.68 0.109 0.051 1.32 15m
MonST3R [77] - - - 0.078 0.038 0.49 1.0s
@@
CasualSAM [79] 0.185 0.022 0.23 0.209 0.027 0.28 2.8m
LEAP-VO [6] 0.167 0.011 0.09 - - - 0.8s
ACE-Zero [3] 0.062 0.012 0.11 0.056 0.012 0.12 1.6s
Particle-SfM [80] 0.081 0.014 0.20 0.087 0.015 0.29 35s
RoDynRF [35] 0.548 0.074 0.70 0.562 0.087 0.90 6.6m
MonST3R [77] - - - 0.690 0.078 0.54 1.0s
@@
Implementation details. In our two-stage training scheme, we first pretrain our model on synthetic data of static scenes, which include 163 scenes from TartanAir [69] and 5K videos from static Kubric [15]. In the second stage, we finetune motion module $F_{m}$ on 11K dynamic videos from Kubric [15]. Each training example consists of a 7-frame video sequence. We set $w_{\text{flow}}=0.02,w_{\text{motion}}=0.1$ during training. Training the camera tracking module using the Adam optimizer [24] takes around 4 days with 8 Nvidia 80G A100s. Within the initialization and frontend phase, we set mono-depth regularization weight $w_{d}=0.05$. In the backend phase, we set $\gamma_{d}=1\times 10^{-4},\beta_{d}=0.05,\tau_{f}=50$, In term of consistent video depth optimization, we set $w_{\text{flow}}=w_{\text{prior}}=1.0,w_{\text{temp}}=0.2$. The average running time of our optimization is 1.3 FPS for video depths at resolution $336\times 144$ on Sintel, but we visualize and evaluate them at resolution of $672\times 288$. We refer readers to supplemental material for more details of network architectures, and other training/inference settings.
Baseline. We compare MegaSaM to recent camera pose estimation methods on both calibrated (known focal length) and uncalibrated (unknown focal length) videos. ACE-Zero [3] is a state-of-the art camera localization method based on scene coordinate regression, designed for static scenes. CasualSAM [79] and RoDynRF [35] jointly estimate camera parameters and dense scene geometry through optimizing mono-depth networks or instant-NGP [37]. Particle-SfM [80] and LEAP-VO [6] estimate cameras from dynamic videos by predicting motion segmentation from long-term trajectories, then using them to mask out moving objects within a standard visual odometry or SfM pipeline. The concurrent work MonST3R [77] extends Dust3R [67] to handle dynamic scenes, estimating camera parameters from global 3D point clouds predicted from pairs of input frames. To evaluate depth accuracy, we compare our outputs to those from CasualSAM, MonST3R, and VideoCrafter [20]. We also include raw mono-depth from DepthAnything-V2 [73] for the sake of completeness. We run all the above baselines using their respective open-source implementations on the same machine with single Nvidia A100 GPU.

### v2

# MegaSaM: Accurate, Fast and Robust Structure and Motion from Casual Dynamic Videos

@@ 1 Introduction
Extracting camera parameters and scene geometry from a set of images is a fundamental problem in computer vision, commonly referred to as structure from motion (SfM) or Simultaneous Localization and Mapping (SLAM). While decades of research have yielded mature algorithms for stationary scenes with large camera baselines, these methods often falter when applied to casual monocular videos captured in uncontrolled setting [kopf2021rcvd, zhang2022structure]. Such videos, frequently captured by handheld cameras, typically exhibit limited camera motion parallax (e.g., nearly stationary or rotational cameras) and a broad range of focal lengths, and often include moving objects and scene dynamics. Recent efforts to address these challenges have focused on two primary strategies: optimizing camera and scene geometry by fine-tuning mono-depth networks or reconstructing radiance fields [luo2020consistent, zhang2021consistent, zhang2022structure, liu2023robust]; or combining intermediate estimates derived from monocular video, such as depth, flows, long-term trajectory and motion segmentation, into a global optimization framework [kopf2021rcvd, chen2024leap, zhang2024monst3r, zhao2022particlesfm, liu2023robust]. However, these approaches are computational expensive or brittle when applied to unconstrained videos that feature long time duration, unconstrained camera paths, or complex scene dynamics.
@@
In this work, we present MegaSaM, a full pipeline for accurate, fast and robust camera tracking and depth estimation from in-the-wild monocular videos of dynamic scenes. Our approach combines the strengths of several prior works, leading to results with previously unachievable quality, as shown in Fig. 1. In particular, we reexamine and extend prior deep visual SLAM framework for camera tracking. A defining feature of deep visual SLAM systems like DROID-SLAM [teed2021droid] is that they adopt a differentiable bundle adjustment (BA) layer that iteratively updates scene geometry and camera pose variables, and intermediate predictions are learned from large amounts of data through camera and flow supervision. We find that such a learned layer is critical for achieving accurate and efficient camera pose estimation for the more challenging case of dynamic video. Building on this foundation, one of our key innovations for handling dynamic scene is to integrate monocular depth priors and motion probability maps into a differentiable SLAM paradigm.
@@ 2 Related Work
SLAM and SfM are used to estimate camera parameters and 3D scene structure from video sequences or unstructured image collections. Conventional approaches tackle this problem by first estimating 2D correspondence between images through feature matching [snavely2006photo, agarwal2011building, schonberger2016structure, pollefeys2004visual, pollefeys2008detailed, campos2021orb, mur2015orb, davison2007monoslam, klein2007parallel, sweeney2019structure] or photometric alignment [newcombe2011dtam, engel2014lsd, engel2017direct]. They then optimize 3D point locations and camera parameters by minimizing reprojection or photo-consistency errors through bundle adjustment (BA) [triggs2000bundle].
Recently, deep visual SLAM and SfM systems have emerged that adopt deep neural networks to estimate pairwise or long-term correspondences [bloesch2018codeslam, tang2018ba, wang2017deepvo, yang2020d3vo, teed2024deep, teed2021droid, wang2024vggsfm, shen2023dytanvo, czarnowski2020deepfactors, he2024detector, kastenfast, holynski2020reducing], to reconstruct radiance fields [lin2021barf, Fu_2024_CVPR, park2023camp] or global 3D point clouds [leroy2024grounding, wang2024dust3r]. While these methods demonstrate accurate camera tracking and reconstruction, they typically assume predominantly static scenes and sufficient camera baselines between frames. Therefore, their performance can degrade significantly or fail entirely in the presence of scene dynamics or limited camera parallax.
Several recent works share similar goals with ours in addressing these limitations. Robust-CVD [kopf2021rcvd] and CasualSAM [zhang2022structure] jointly estimate camera parameters and dense depth maps from dynamic videos by optimizing a spatially varying spline or fine-tuning monocular depth networks. Particle-SfM [zhao2022particlesfm] and LEAP-VO [chen2024leap] first infer moving object masks based on long-range trajectories, then use this information to downweight the contribution of the features during bundle adjustment. Concurrent work, MonST3R [zhang2024monst3r], adopts a 3D point cloud representation from DuST3R [wang2024dust3r] and localizes cameras via an additional alignment optimization. Our approach shares similar ideas, but we show that performance can be significantly improved by coupling the differentiable SLAM system with the intermediate predictions of underlying dynamic scenes.
@@
Recent work on monocular depth prediction has shown strong generalization on in-the-wild single images by training deep neural networks on large amounts of synthetic and real-world data [li2018megadepth, li2019learning, ranftl2021vision, yin2021learning, ranftl2020towards, saxena2024surprising, depthanything, yang2024depth, yin2023metric3d, piccinelli2024unidepth, ke2024repurposing, godard2017unsupervised]. However, these single-image models tend to produce temporally inconsistent depth from videos. To overcome this issue, prior techniques propose to fine-tune mono-depth models by performing test-time optimization [luo2020consistent, zhang2021consistent] or to use transformer or diffusion models to directly predict video depths [wang2023neural, shao2024learning, hu2024depthcrafter]. Our approach follows the spirit of the first paradigm, but we show that we can achieve better video depth quality without resorting to expensive network fine-tuning for every video.
@@
Several recent works have adopted time-varying radiance field representations to perform dynamic scene reconstruction and novel view synthesis from in-the-wild videos [park2021nerfies, park2021hypernerf, li2021neural, li2023dynibar, gao2022monocular, wang2024gflow, lei2024mosca, wang2024shape, liu2023robust, wu2024cat4d]. Our work is orthogonal to most of these techniques since most radiance field reconstruction methods require camera parameters or video depth maps as inputs, and our outputs can be used as inputs to these systems.
@@ 3 MegaSaM
Given an unconstrained, continuous video sequence $\mathcal{V}=\{I_{i}\in\mathcal{R}^{H\times W}\}^{N}_{i=1}$ our goal is to estimate camera poses $\hat{\mathbf{G}}_{i}\in SE(3)$, focal length $\hat{f}$ (if unknown), and dense video depth maps $\mathcal{D}=\{\hat{D}_{i}\}^{N}_{i=1}$. Our approach does not pose any constraints on the camera and object motions present in the input video. Our camera tracking and video depth estimation modules build upon prior deep visual SLAM (in particular, DROID-SLAM [teed2021droid]) and casual structure and motion [zhang2022structure] frameworks, respectively.
@@ 3.1 Deep visual SLAM formulation
Deep visual SLAM systems like DROID-SLAM [teed2021droid] are characterized by a differentiable, learned bundle adjustment (BA) layer that iteratively updates structure and motion parameters. In particular, they keep track of two state variables while processing a video: a per-frame low-resolution disparity map $\hat{\mathbf{d}}_{i}\in\mathcal{R}^{\frac{H}{8}\times\frac{W}{8}}$, and camera poses $\hat{\mathbf{G}}_{i}\in SE(3)$. These variables are updated iteratively during both training and inference stages through the differentiable BA layer, which operates over a set of image pairs from a frame-graph $(I_{i},I_{j})\in\mathcal{P}$, built dynamically to connect frames with overlapping field of view.
@@
Differentiable bundle adjustment. DROID-SLAM assumes known focal length, but focal length is typically not known a priori for in-the-wild videos. Therefore, we optimize camera poses, focal length and disparity by iteratively minimizing a weighted reprojection cost between the current flows predicted by the network and the rigid-motion ones derived from camera parameters and disparity [hagemann2023deep]:
@@
[…]
@@
[…]
@@
[…]
@@ 3.2.1 Training
Disparity and camera initialization. DROID-SLAM initializes disparity $\hat{\mathbf{d}}$ by simply setting it to a constant value of $1$. However, we find that this initialization fails to perform accurate camera tracking on videos with limited camera baselines and complex scene dynamics. Inspired by recent work [li2021neural, wang2024shape, li2023dynibar, liu2023robust], during both the training and inference stages, we perform data-driven initialization by integrating a mono-depth prior. During training, we initialize $\hat{\mathbf{d}}$ with disparity from DepthAnything [depthanything] with the estimated global scale and shift borrowed from the ground truth depth of each training sequence. For each training sequence, we initialize first two camera poses to the ground-truth to remove the gauge ambiguity, and initialize the camera focal length by randomly perturbing the ground truth value by $25\%$.
Refer to caption Refer to caption
Refer to caption Refer to caption
I Image I Motion probability
@@ 3.2.2 Inference
Initialization and frontend tracking. Similar to training, we incorporate both mono-depth and focal length predictions into the inference pipeline. In particular, we initialize per-frame disparity maps $\hat{\mathbf{d}}_{i}$ with metric aligned monocular disparity $D^{\text{align}}_{i}=\hat{\alpha}D^{\text{rel}}_{i}+\hat{\beta}$, where $D^{\text{rel}}_{i}$ is per-frame affine-invariant disparity from [depthanything], and per-video global scale and shift parameters $(\hat{\alpha},\hat{\beta})$ are estimated through median aligning $D^{\text{rel}}_{i}$ with an additional metric depth estimates $D^{\text{abs}}_{i}$ from UniDepth [piccinelli2024unidepth]: $\hat{\alpha}_{i}=\frac{D^{\text{abs}}_{i}-\text{median}_{i}(D^{\text{abs}}_{i})}{D^{\text{rel}}_{i}-\text{median}(D^{\text{rel}}_{i})}$; $\hat{\beta}=\text{median}\left(D^{\text{abs}}_{i}-\hat{\alpha}D^{\text{rel}}_{i}\right)$. [^1] We combine two mono-depth models since they have complementary advantages: DepthAnything provides more accurate and consistent depths, whereas UniDepth provides scene scale and shift estimates. The UniDepth model also predicts the focal length of each frame; we use the median estimate across the video frames to obtain an initial focal length estimate $\hat{f}$, which is fixed within the frontend stage.
@@
To see why, we explore the approximate Hessian matrix from the linear system in Eq. 4. As shown by Goli et al. [goli2024bayes], given the posterior, $p(\theta|\mathcal{I})$, we can use Laplace approximations to estimate the covariance $\Sigma$ of the variables through the inverse Hessian: $\Sigma_{\theta}=-\mathbf{H}(\theta^{*})^{-1}$, where $\theta^{*}$ is the MAP estimate of the parameters and $\Sigma_{\theta}$ denotes the epistemic uncertainty of estimated variables [kendall2017uncertainties]. Since inverting the full Hessian is computationally expensive when the number of input frames is large, we follow Ritter et al. [ritter2018scalable] and approximate $\Sigma_{\theta}$ through the diagonal of the Hessian:
@@ 3.3 Consistent depth optimization
In particular, we follow CasualSAM [zhang2022structure] and perform an additional first-order optimization on video depths along with per-frame aleatoric uncertainty maps. Our objective consist of three cost functions:
@@
where $\mathcal{C}_{\text{flow}}$ denotes pairwise 2D flow reprojection loss, $\mathcal{C}_{\text{temp}}$ is temporal depth consistency loss, and $\mathcal{C}_{\text{prior}}$ is scale invariant mono-depth prior loss. We derive 2D optical flow at the original frame resolution from an off-the-shelf module [teed2020raft].
Note that our design has a few differences compared to CasualSAM: (i) instead of performing a time-consuming mono-depth network fine-tuning, we construct and optimize a sequence of variables for disparity and uncertainty over the input video; (ii) we fix camera parameters instead of jointly optimizing cameras and depths during optimization; (iii) we adopt surface normal consistency and multi-scale depth gradient matching losses [li2018megadepth, sayed2022simplerecon] to replace the depth prior loss used in CasualSAM [zhang2022structure]. We find that these modifications lead to much faster optimization times as well as more accurate video depth estimates. We refer readers see the supplementary material for more details of our losses and optimization scheme.
@@ 4 Experiments
CasualSAM [zhang2022structure] 0.036 0.013 0.20 0.067 0.019 0.47 1.6m
LEAP-VO [chen2024leap] 0.041 0.023 0.17 - - - 1.3s
ACE-Zero [brachmann2024scene] 0.053 0.028 0.30 0.065 0.028 1.92 10s
Particle-SfM [zhao2022particlesfm] 0.062 0.032 1.26 0.057 0.038 1.64 21s
RoDynRF [liu2023robust] 0.110 0.049 1.68 0.109 0.051 1.32 15m
MonST3R [zhang2024monst3r] - - - 0.078 0.038 0.49 1.0s
@@
CasualSAM [zhang2022structure] 0.185 0.022 0.23 0.209 0.027 0.28 2.8m
LEAP-VO [chen2024leap] 0.167 0.011 0.09 - - - 0.8s
ACE-Zero [brachmann2024scene] 0.062 0.012 0.11 0.056 0.012 0.12 1.6s
Particle-SfM [zhao2022particlesfm] 0.081 0.014 0.20 0.087 0.015 0.29 35s
RoDynRF [liu2023robust] 0.548 0.074 0.70 0.562 0.087 0.90 6.6m
MonST3R [zhang2024monst3r] - - - 0.690 0.078 0.54 1.0s
@@
Implementation details. In our two-stage training scheme, we first pretrain our model on synthetic data of static scenes, which include 163 scenes from TartanAir [tartanair2020iros] and 5K videos from static Kubric [greff2021kubric]. In the second stage, we finetune motion module $F_{m}$ on 11K dynamic videos from Kubric [greff2021kubric]. Each training example consists of a 7-frame video sequence. We set $w_{\text{flow}}=0.02,w_{\text{motion}}=0.1$ during training. Training the camera tracking module using the Adam optimizer [Kingma2014AdamAM] takes around 4 days with 8 Nvidia 80G A100s. Within the initialization and frontend phase, we set mono-depth regularization weight $w_{d}=0.05$. In the backend phase, we set $\gamma_{d}=1\times 10^{-4},\beta_{d}=0.05,\tau_{f}=50$, In term of consistent video depth optimization, we set $w_{\text{flow}}=w_{\text{prior}}=1.0,w_{\text{temp}}=0.2$. The average running time of our optimization is 1.3 FPS for video depths at resolution $336\times 144$ on Sintel, but we visualize and evaluate them at resolution of $672\times 288$. We refer readers to supplemental material for more details of network architectures, and other training/inference settings.
Baseline. We compare MegaSaM to recent camera pose estimation methods on both calibrated (known focal length) and uncalibrated (unknown focal length) videos. ACE-Zero [brachmann2024scene] is a state-of-the art camera localization method based on scene coordinate regression, designed for static scenes. CasualSAM [zhang2022structure] and RoDynRF [liu2023robust] jointly estimate camera parameters and dense scene geometry through optimizing mono-depth networks or instant-NGP [mueller2022instant]. Particle-SfM [zhao2022particlesfm] and LEAP-VO [chen2024leap] estimate cameras from dynamic videos by predicting motion segmentation from long-term trajectories, then using them to mask out moving objects within a standard visual odometry or SfM pipeline. The concurrent work MonST3R [zhang2024monst3r] extends Dust3R [wang2024dust3r] to handle dynamic scenes, estimating camera parameters from global 3D point clouds predicted from pairs of input frames. To evaluate depth accuracy, we compare our outputs to those from CasualSAM, MonST3R, and VideoCrafter [hu2024depthcrafter]. We also include raw mono-depth from DepthAnything-V2 [yang2024depth] for the sake of completeness. We run all the above baselines using their respective open-source implementations on the same machine with single Nvidia A100 GPU.